# arXiv:2512.09295v2[math.ST] 01 Mar 2026

### v2

@@ 3.1 G-expansion and Denoising Equations
Theorem 3 (G-expansion of the Optimal Transport Map).
Consider the same setting as Theorem 1, we define the following series expansion
@@
where $h_{1},h_{2},\ldots$ are functions defined iteratively through the Bell polynomials $B_{n,k}(h_{1},h_{2},\ldots,h_{n-k+1})$ in Definition 3:
@@
Then, for $y\in\mathbb{R}$ such that $T_{G}(y)$ is absolutely convergent, $T_{G}$ is the optimal transport map pushing $Q$ onto $P$ as in Definition 2, namely,
@@
Remark 4.
The crucial difference between Theorem 3 and Theorem 1 is that the expansion in Theorem 3 depends on the higher-order score functions of $Q$, which is identifiable and estimable from the noisy measurements $Y$. This makes Theorem 3 practical for denoising applications. We evaluate a few initial terms explicitly before presenting the proof.
@@
    $B_{4,1}(h_{1},h_{2},h_{3},h_{4})=h_{4},\quad B_{4,2}(h_{1},h_{2},h_{3})=4h_{1}h_{3}+3h_{2}^{2},\quad B_{4,3}(h_{1},h_{2})=6h_{1}^{2}h_{2},\quad B_{4,4}(h_{1})=h_{1}^{4}.$
    Equation (3.2) reads $G^{(1)}B_{4,1}(h_{1},h_{2},h_{3},h_{4})+G^{(2)}B_{4,2}(h_{1},h_{2},h_{3})+G^{(3)}B_{4,3}(h_{1},h_{2})+G^{(4)}B_{4,4}(h_{1})-4G^{(3)}B_{3,1}(h_{1},h_{2},h_{3})-4G^{(4)}B_{3,2}(h_{1},h_{2})-4G^{(5)}B_{3,3}(h_{1})+6G^{(5)}B_{2,1}(h_{1},h_{2})+6G^{(6)}B_{2,2}(h_{1})-4G^{(7)}B_{1,1}(h_{1})+G^{(8)}=0$, and thus we have
@@
    $-\frac{G^{(4)}}{G^{(1)}}h_{1}^{4}-6\frac{G^{(3)}}{G^{(1)}}h_{1}^{2}h_{2}-\frac{G^{(2)}}{G^{(1)}}(4h_{1}h_{3}+3h_{2}^{2})\;.$
Proof of Theorem 3.
@@
Recall if the series expansion of $T_{G}(y)=y+\sum_{k=1}^{\infty}\frac{\eta^{k}}{k!}h_{k}(y)$ is absolutely convergent, we can apply Lemma 2 that expands $F$ in terms of derivatives of $G$, and then by Fubini’s theorem, we have
@@
Apply Lemma 3 that relates powers of series to Bell polynomials, we have again by Fubini’s theorem,
@@
Following the arguments in Theorem 3 and Lemma 5, we have the following corollary on the finite-order denoising accuracy, measured in terms of the integral form of the static Monge-Ampère equation [Caf92, Vil08, DL25, Lia25], namely $|F\circ T_{K}(y)-G(y)|$ at any given $y\in\mathbb{R}$. The proof is deferred to Section A.3. As noted earlier, we can assume that $\sigma\in[0,1]$ without loss of generality since linear rescaling of $Y$ does not change the denoising problem. Therefore, we can always assume $\eta=\sigma^{2}/2\in[0,1/2]$.
Corollary 1.
Consider the same setting as Theorem 3 and $\{h_{k}\}_{k\geq 1}$ defined therein. For any integer $K\geq 1$, define the $K$-th order denoiser as
@@
Theorem 3 and Corollary 1 provide an iterative refinement scheme to construct higher-order denoisers based on polynomials of $\tfrac{G^{(k+1)}}{G^{(1)}},k\geq 1$. Since $\tfrac{G^{(k+1)}}{G^{(1)}}=\tfrac{q^{(k)}}{q}$ is the higher-order score function of $Q$, it is identifiable from the noisy observations $\{Y_{i}\}_{i=1}^{n}$ drawn i.i.d. from $Q$.
In the hierarchy of denoisers $\{T_{K}\}_{K\geq 1}$ defined above, we start with the trivial denoiser $T_{0}(y)=y$, then use the first-order score function of $G$ to refine $T_{1}(y)=y+\eta\frac{G^{(2)}(y)}{G^{(1)}(y)}$. Theorem 3 generalizes this procedure to arbitrary $K$-order, with each refinement step involving a polynomial of the higher-order score functions of $\frac{G^{(m)}(y)}{G^{(1)}(y)}$ up to order $m\leq 2K$. The recursions are defined using Bell-type polynomials in (3.2). The denoising accuracy improves as $\eta^{K}$, as quantified in Corollary 1 (and Theorem 2). In the limit $K\rightarrow\infty$, this denoiser recovers the optimal transport map $T_{K}\rightarrow F^{-1}\circ G$, without requiring the knowledge of $F$, the signal distribution.
@@ 3.2 Plug-in Estimation via Gaussian Kernel Smoothing
where $h_{1},h_{2},\ldots,h_{K}$ are defined by the Bell polynomial type recursion in Theorem 3 that involve only the derivatives of $\{G^{(m)}\}_{m=1}^{2K}$. Therefore, to estimate $T_{K}$, it suffices to estimate the density function $q$ and its derivatives up to order 2K12K-1.
Given i.i.d. samples from $Q$, denoted as $\{Y_{i}\}_{i=1}^{n}$, we can estimate $q$ and its derivatives using Gaussian kernel smoothing. Specify a bandwidth $b>0$, we define the following estimator for $q^{(m)}$, the $m$-th derivative of $q$:
@@
Theorem 4 (Estimation Rates for $q^{(m)}$: Gaussian Kernel Smoothing).
@@
where the last line uses Lemma 7. With the choice $b=b_{m}=\left(\sqrt{\tfrac{8}{\pi}}\tfrac{m!}{L}\right)^{\frac{1}{2m+5}}\cdot n^{-\frac{1}{2m+5}}$, we finish the proof. ∎
Remark 5.
One can rationalize the estimation rate as follows: the optimal rate for estimating the $m$-th derivative of a density function of smoothness $\alpha=m+2$ is of the rate
@@ 3.3 Direct Estimation via Higher-Order Score Matching
Theorem 5 (Estimation Rates for $q^{(m)}/q$: Higher-Order Score Matching).
@@
Proof of Theorem 5.
We first apply the representation lemma for higher-order score functions (see Lemma 8) with $f=\widehat{f}_{m}$,
$\mathop{\mathbb{E}}_{Y\sim Q}\frac{1}{2}\left|\widehat{f}_{m}(Y)-f^{\ast}_{m}(Y)\right|^{2}$ $=\mathop{\mathbb{E}}_{Y\sim Q}\left[\frac{1}{2}\widehat{f}_{m}(Y)^{2}+(-1)^{m+1}\widehat{f}_{m}^{(m)}(Y)\right]-\mathop{\mathbb{E}}_{Y\sim Q}\left[\frac{1}{2}f^{\ast}_{m}(Y)^{2}+(-1)^{m+1}(f^{\ast}_{m})^{(m)}(Y)\right]\;.$ (3.5)
@@
Due to Giné-Zinn symmetrization (see Lemma 9), we have
@@
[…]
@@ Acknowledgements
The author acknowledges the support from the NSF Career Award (DMS-2042473) and the Wallman Society of Fellows at the University of Chicago. The author also thanks Nikos Ignatiadis, Ben Recht, and Richard Samworth for comments on an earlier draft. This work is part II of a two-part series on distributional shrinkage; part I can be found in [Lia25].
@@ Appendix A Technical Proofs
Lemma 1.
Consider the model in Definition 1 with an analytic $F$, we have $G(y)=\sum_{k=0}^{\infty}\frac{\eta^{k}}{k!}F^{(2k)}(y)\;.$
Proof of Lemma 1.
[…]
@@
Lemma 2.
Consider the model in Definition 1 with an analytic $F$, we have $F(y)=\sum_{k=0}^{\infty}\frac{(-\eta)^{k}}{k!}G^{(2k)}(y)\;.$
Proof of Lemma 2.
The proof proceeds by recalling Lemma 1, namely $G^{(2k)}(y)=\sum_{l=0}^{\infty}\frac{\eta^{l}}{l!}F^{(2k+2l)}(y)$. We have
@@
Lemma 3.
Recall the definition of Bell polynomials $B_{n,j}(x_{1},x_{2},\ldots,x_{n-j+1})$ in Definition 3, we have
@@
Proof of Lemma 3.
$\frac{1}{j!}\left(\sum_{k=1}^{\infty}\frac{\eta^{k}}{k!}x_{k}\right)^{j}$ $=\frac{1}{j!}\sum_{i_{1}+i_{2}+\cdots=j}\frac{j!}{i_{1}!i_{2}!\cdots i_{n-j+1}!}\prod_{m=1}^{n-j+1}\left(\frac{\eta^{m}}{m!}x_{m}\right)^{i_{m}}$
@@
Now we are ready to prove Theorem 1.
Proof of Theorem 1.
@@
With the definition of $g_{1},g_{2},\ldots$ in (1), and the fact $B_{k,0}=\mathbb{I}(k=0)$, we have
$\sum_{j=0}^{k}F^{(j)}(y)\cdot B_{k,j}\big(g_{1},\dots,g_{k-j+1}\big)(y)=\begin{cases}F^{(0)}(y)\;,&k=0\\ F^{(1)}(y)B_{1,1}(g_{1})=F^{(2)}(y)\;,&k=1\\ \sum_{j=1}^{k}F^{(j)}(y)\cdot B_{k,j}\big(g_{1},\dots,g_{k-j+1}\big)(y)=F^{(2k)}(y)\;,&k\geq 2\end{cases}\;,$
@@
Here, the last line uses Lemma 1. By the uniqueness of the optimal transport map, we conclude that $F^{-1}\circ G(y)=T_{F}(y)$ for all $y$ such that the series expansion is absolutely convergent. ∎
@@
[…]
Proof of Lemma 4.
@@
[…]
@@
Lemma 5.
Under the same setting as in Theorem 2,
@@
Proof of Lemma 5.
First, recall Lemma 1,
@@
Here the constant $C_{2}>0$ depends on the upper bounds of $\sup_{y\in\mathrm{supp}(P)}|F^{(K+1)}(y)|$ and $\sup_{y\in\mathrm{supp}(P)}|g_{k}(y)|$ for $k=1,2,\ldots,K$. Recall the definition of $g_{k}(y)$ in Theorem 2, we know that these are finite-degree polynomials of $\{F^{(k)}/F^{(1)}\}_{k=2}^{2K}$. Under the assumption that $\sup_{y\in\mathrm{supp}(P)}|F^{(k)}(y)|$ and $\inf_{y\in\mathrm{supp}(P)}|F^{(1)}(y)|$ bounded, we have $g_{k}$’s uniformly bounded.
@@
$=\frac{\eta^{k}}{k!}\sum_{j=0}^{k}F^{(j)}(y)\cdot B_{k,j}\big(g_{1},\dots,g_{k-j+1}\big)(y)=\frac{\eta^{k}}{k!}F^{(2k)}\;.$
@@
$\left|F\circ T_{K}(y)-G(y)\right|$ $\leq\left|F\circ T_{K}(y)-\sum_{j=0}^{K}F^{(j)}(y)\frac{1}{j!}\left(\sum_{k=1}^{K}\frac{\eta^{k}}{k!}g_{k}(y)\right)^{j}\right|+\left|G(y)-\sum_{k=0}^{K}\frac{\eta^{k}}{k!}F^{(2k)}(y)\right|$
$\qquad\qquad+\left|\sum_{j=0}^{K}F^{(j)}(y)\frac{1}{j!}\left(\sum_{k=1}^{K}\frac{\eta^{k}}{k!}g_{k}(y)\right)^{j}-\sum_{k=0}^{K}\frac{\eta^{k}}{k!}F^{(2k)}(y)\right|$
@@
Lemma 6.
Consider the model in Definition 1 and assume $F(y)$ has derivatives up to order $2K+2$ uniformly bounded over $y\in\mathbb{R}$. Then
@@
Proof of Lemma 6.
Note from Lemma 1,
@@
Proof of Corollary 1.
First, we note that by Lemma 6, if $F^{(k)}$ are bounded for $k\leq 2K+2$, then
@@
Following the logic behind the calculations in Theorem 3, we have
@@
$=\sum_{k=0}^{K}\frac{\eta^{k}}{k!}\sum_{l=0}^{k}(-1)^{l}\frac{k!}{l!(k-l)!}\sum_{j=0}^{k-l}G^{(2l+j)}(y)B_{k-l,j}\big(h_{1},\dots,h_{k-l-j+1}\big)(y)+O(\eta^{K+1})\;.$
@@
Lemma 7 (Bias and Variance of Gaussian Smoothing).
@@
[…]
@@
[…]
@@
[…]
@@
Lemma 8 (Higher-order Score Representation).
@@
$\mathop{\mathbb{E}}_{Y\sim Q}\frac{1}{2}\left|f(Y)-f^{\ast}_{m}(Y)\right|^{2}=\mathop{\mathbb{E}}_{Y\sim Q}\left[\frac{1}{2}f(Y)^{2}+(-1)^{m+1}f^{(m)}(Y)\right]-\mathop{\mathbb{E}}_{Y\sim Q}\left[\frac{1}{2}f^{\ast}_{m}(Y)^{2}+(-1)^{m+1}(f^{\ast}_{m})^{(m)}(Y)\right]\;.$
Proof of Lemma 8.
@@
[…]
Here, the last step applies integration by parts $m$ times and uses the vanishing boundary conditions up to order m1m-1. Plug in $f=f^{\ast}_{m}$, we have the higher-order Fisher-type information [Bob24] identity
[…]

### v3

@@ 3.1 G-expansion and Denoising Equations
Theorem 3.2 (G-expansion of the Optimal Transport Map).
Consider the same setting as Theorem 2.4, we define the following series expansion
@@
where $h_{1},h_{2},\ldots$ are functions defined iteratively through the Bell polynomials $B_{n,k}(h_{1},h_{2},\ldots,h_{n-k+1})$ in Definition 2.3:
@@
Then, for $y\in\mathbb{R}$ such that $T_{G}(y)$ is absolutely convergent, $T_{G}$ is the optimal transport map pushing $Q$ onto $P$ as in Definition 2.2, namely,
@@
Remark 3.3.
The crucial difference between Theorem 3.2 and Theorem 2.4 is that the expansion in Theorem 3.2 depends on the higher-order score functions of $Q$, which are identifiable and estimable from the noisy measurements $Y$. This makes Theorem 3.2 practical for denoising applications. We evaluate a few initial terms explicitly before presenting the proof.
@@
    $B_{4,1}=h_{4},\quad B_{4,2}=4h_{1}h_{3}+3h_{2}^{2},\quad B_{4,3}=6h_{1}^{2}h_{2},\quad B_{4,4}=h_{1}^{4}.$
    Substituting into Equation (3.2) and simplifying, we have
@@
Proof of Theorem 3.2.
@@
Recall if the series expansion of $T_{G}(y)=y+\sum_{k=1}^{\infty}\frac{\eta^{k}}{k!}h_{k}(y)$ is absolutely convergent, we can apply Lemma A.2 that expands $F$ in terms of derivatives of $G$, and then by Fubini’s theorem, we have
@@
Apply Lemma A.3 that relates powers of series to Bell polynomials, we have again by Fubini’s theorem,
@@
Following the arguments in Theorem 3.2 and Lemma A.5, we have the following corollary on the finite-order denoising accuracy, measured in terms of the integral form of the static Monge-Ampère equation [Caf92, Vil08, DL25, Lia25], namely $|F\circ T_{K}(y)-G(y)|$ at any given $y\in\mathbb{R}$. The proof is deferred to Section A.3. As noted earlier, we can assume that $\sigma\in[0,1]$ without loss of generality since linear rescaling of $Y$ does not change the denoising problem. Therefore, we can always assume $\eta=\sigma^{2}/2\in[0,1/2]$.
Corollary 3.4.
Consider the same setting as Theorem 3.2 and $\{h_{k}\}_{k\geq 1}$ defined therein. For any integer $K\geq 1$, define the $K$-th order denoiser as
@@
Theorem 3.2 and Corollary 3.4 provide an iterative refinement scheme to construct higher-order denoisers based on polynomials of $\tfrac{G^{(k+1)}}{G^{(1)}},k\geq 1$. Since $\tfrac{G^{(k+1)}}{G^{(1)}}=\tfrac{q^{(k)}}{q}$ is the higher-order score function of $Q$, it is identifiable from the noisy observations $\{Y_{i}\}_{i=1}^{n}$ drawn i.i.d. from $Q$.
In the hierarchy of denoisers $\{T_{K}\}_{K\geq 1}$ defined above, we start with the trivial denoiser $T_{0}(y)=y$, then use the first-order score function of $G$ to refine $T_{1}(y)=y+\eta\frac{G^{(2)}(y)}{G^{(1)}(y)}$. Theorem 3.2 generalizes this procedure to arbitrary order $K$, with each refinement step involving a polynomial in the higher-order score functions $\frac{G^{(m)}(y)}{G^{(1)}(y)}$ up to order $m\leq 2K$. The recursions are defined using Bell-type polynomials in (3.2). The denoising accuracy improves at rate $\eta^{K}$, as quantified in Corollary 3.4 (and Theorem 2.6). In the limit $K\rightarrow\infty$, this denoiser recovers the optimal transport map $T_{K}\rightarrow F^{-1}\circ G$, without requiring the knowledge of $F$, the signal distribution.
@@ 3.2 Plug-in Estimation via Gaussian Kernel Smoothing
where $h_{1},h_{2},\ldots,h_{K}$ are defined by the Bell-polynomial-type recursion in Theorem 3.2 that involves only the derivatives of $\{G^{(m)}\}_{m=1}^{2K}$. Therefore, to estimate $T_{K}$, it suffices to estimate the density function $q$ and its derivatives up to order 2K12K-1.
Given i.i.d. samples from $Q$, denoted as $\{Y_{i}\}_{i=1}^{n}$, we can estimate $q$ and its derivatives using Gaussian kernel smoothing. Given a bandwidth $b>0$, we define the following estimator for $q^{(m)}$, the $m$-th derivative of $q$:
@@
Theorem 3.5 (Estimation Rates for $q^{(m)}$: Gaussian Kernel Smoothing).
@@
where the last line uses Lemma A.7. With the choice $b=b_{m}=\left(\sqrt{\tfrac{8}{\pi}}\tfrac{m!}{L}\right)^{\frac{1}{2m+5}}\cdot n^{-\frac{1}{2m+5}}$, we finish the proof. ∎
Remark 3.6.
One can rationalize the estimation rate as follows: the optimal rate for estimating the $m$-th derivative of a density function of smoothness $\alpha=m+2$ is at the rate
@@ 3.3 Direct Estimation via Higher-Order Score Matching
Theorem 3.7 (Estimation Rates for $q^{(m)}/q$: Higher-Order Score Matching).
@@
Proof of Theorem 3.7.
We first apply the representation lemma for higher-order score functions (see Lemma A.8) with $f=\widehat{f}_{m}$,
$\mathop{\mathbb{E}}_{Y\sim Q}\frac{1}{2}\left|\widehat{f}_{m}(Y)-f^{\ast}_{m}(Y)\right|^{2}$ $=\mathop{\mathbb{E}}_{Y\sim Q}\left[\tfrac{1}{2}\widehat{f}_{m}(Y)^{2}+(-1)^{m+1}\widehat{f}_{m}^{(m)}(Y)\right]$
$\qquad-\mathop{\mathbb{E}}_{Y\sim Q}\left[\tfrac{1}{2}f^{\ast}_{m}(Y)^{2}+(-1)^{m+1}(f^{\ast}_{m})^{(m)}(Y)\right]\;.$ (3.5)
@@
Due to Giné-Zinn symmetrization (see Lemma A.9), we have
@@
$\mathop{\mathbb{E}}_{\{\epsilon_{i}\}_{i=1}^{n}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(Y_{i})$
$\leq\inf_{0\leq\delta\leq L}\left(4\delta+\frac{12}{\sqrt{n}}\int_{\delta}^{L}\sqrt{\log\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{\infty})}\,\mathrm{d}\epsilon\right)\leq\begin{cases}C_{1}n^{-\alpha},&\alpha<\frac{1}{2}\\ C_{2}n^{-\frac{1}{2}}\log n,&\alpha=\frac{1}{2}\\ C_{3}n^{-\frac{1}{2}},&\alpha>\frac{1}{2}\end{cases}$
@@ Acknowledgements
The author acknowledges the support from the NSF Career Award (DMS-2042473) and the Wallman Society of Fellows at the University of Chicago. The author also thanks Nabarun Deb, Nikos Ignatiadis, Ben Recht, and Richard Samworth for comments on an earlier draft. This work is part II of a two-part series on distributional shrinkage; part I can be found in [Lia25].
@@ Appendix A Technical Proofs
Lemma A.1.
Consider the model in Definition 2.1 with an analytic $F$, we have $G(y)=\sum_{k=0}^{\infty}\frac{\eta^{k}}{k!}F^{(2k)}(y)\;.$
Proof of Lemma A.1.
Recall that $G(y)=\int F(y-\sigma z)\phi(z)\,\mathrm{d}z$, where $\phi(z)$ is the standard normal density function. Therefore,
@@
Lemma A.2.
Consider the model in Definition 2.1 with an analytic $F$, we have $F(y)=\sum_{k=0}^{\infty}\frac{(-\eta)^{k}}{k!}G^{(2k)}(y)\;.$
Proof of Lemma A.2.
The proof proceeds by recalling Lemma A.1, namely $G^{(2k)}(y)=\sum_{l=0}^{\infty}\frac{\eta^{l}}{l!}F^{(2k+2l)}(y)$. We have
@@
Lemma A.3.
Recall the definition of Bell polynomials $B_{n,j}(x_{1},x_{2},\ldots,x_{n-j+1})$ in Definition 2.3, we have
@@
Proof of Lemma A.3.
$\frac{1}{j!}\left(\sum_{k=1}^{\infty}\frac{\eta^{k}}{k!}x_{k}\right)^{j}$ $=\frac{1}{j!}\sum_{i_{1}+i_{2}+\cdots=j}\frac{j!}{i_{1}!i_{2}!\cdots}\prod_{m=1}^{\infty}\left(\frac{\eta^{m}}{m!}x_{m}\right)^{i_{m}}$
@@
Now we are ready to prove Theorem 2.4.
Proof of Theorem 2.4.
@@
With the definition of $g_{1},g_{2},\ldots$ in (2.4), and the fact $B_{k,0}=\mathbb{I}(k=0)$, we have
$\sum_{j=0}^{k}F^{(j)}(y)\cdot B_{k,j}\big(g_{1},\dots,g_{k-j+1}\big)(y)=\begin{cases}F^{(0)}(y)\;,&k=0\\ F^{(1)}(y)B_{1,1}(g_{1})=F^{(2)}(y)\;,&k=1\\ \sum_{j=1}^{k}F^{(j)}B_{k,j}\big(g_{1},\dots,g_{k-j+1}\big)=F^{(2k)}\;,&k\geq 2\end{cases}\;,$
@@
Here, the last line uses Lemma A.1. By the uniqueness of the optimal transport map, we conclude that $F^{-1}\circ G(y)=T_{F}(y)$ for all $y$ such that the series expansion is absolutely convergent. ∎
@@
$W_{r}^{r}(T_{0}\sharp Q,P)=\int_{\mathbb{R}}|T_{0}(y)-F^{-1}\circ G(y)|^{r}q(y)\,\mathrm{d}y\;.$
Proof of Lemma A.4.
@@
$W_{r}^{r}(T_{0}\sharp Q,P)$ $=\int_{0}^{1}|F_{T_{0}\sharp Q}^{-1}(t)-F^{-1}_{P}(t)|^{r}\,\mathrm{d}t=\int_{0}^{1}|T_{0}\circ G^{-1}(t)-F^{-1}(t)|^{r}\,\mathrm{d}t$
$=\int_{\mathbb{R}}|T_{0}(y)-F^{-1}\circ G(y)|^{r}q(y)\,\mathrm{d}y\;.$
@@
Lemma A.5.
Under the same setting as in Theorem 2.6,
@@
Proof of Lemma A.5.
First, recall Lemma A.1,
@@
Here the constant $C_{2}>0$ depends on the upper bounds of $\sup_{y\in\mathrm{supp}(P)}|F^{(K+1)}(y)|$ and $\sup_{y\in\mathrm{supp}(P)}|g_{k}(y)|$ for $k=1,2,\ldots,K$. Recall the definition of $g_{k}(y)$ in Theorem 2.6, we know that these are finite-degree polynomials of $\{F^{(k)}/F^{(1)}\}_{k=2}^{2K}$. Under the assumption that $\sup_{y\in\mathrm{supp}(P)}|F^{(k)}(y)|$ and $\inf_{y\in\mathrm{supp}(P)}|F^{(1)}(y)|$ bounded, we have $g_{k}$’s uniformly bounded.
@@
$=\frac{\eta^{k}}{k!}\sum_{j=0}^{k}F^{(j)}(y)\cdot B_{k,j}\big(g_{1},\dots,g_{k-j+1}\big)(y)=\frac{\eta^{k}}{k!}F^{(2k)}(y)\;.$
@@
$\left|F\circ T_{K}(y)-G(y)\right|$ $\leq\underbrace{\left|F\circ T_{K}(y)-\sum_{j=0}^{K}\frac{F^{(j)}(y)}{j!}\left(\sum_{k=1}^{K}\frac{\eta^{k}}{k!}g_{k}(y)\right)^{\!j}\right|}_{C_{1}\cdot\eta^{K+1}}$
$\qquad+\underbrace{\left|G(y)-\sum_{k=0}^{K}\frac{\eta^{k}}{k!}F^{(2k)}(y)\right|}_{C_{2}\cdot\eta^{K+1}}$
$\qquad+\underbrace{\left|\sum_{j=0}^{K}\frac{F^{(j)}(y)}{j!}\left(\sum_{k=1}^{K}\frac{\eta^{k}}{k!}g_{k}(y)\right)^{\!j}-\sum_{k=0}^{K}\frac{\eta^{k}}{k!}F^{(2k)}(y)\right|}_{C_{3}\cdot\eta^{K+1}}$
@@
Lemma A.6.
Consider the model in Definition 2.1 and assume $F(y)$ has derivatives up to order $2K+2$ uniformly bounded over $y\in\mathbb{R}$. Then
@@
Proof of Lemma A.6.
Note from Lemma A.1,
@@
Proof of Corollary 3.4.
First, we note that by Lemma A.6, if $F^{(k)}$ are bounded for $k\leq 2K+2$, then
@@
Following the logic behind the calculations in Theorem 3.2, we have
@@
$=\sum_{k=0}^{K}\frac{\eta^{k}}{k!}\sum_{l=0}^{k}(-1)^{l}\binom{k}{l}\sum_{j=0}^{k-l}G^{(2l+j)}(y)B_{k-l,j}\big(h_{1},\dots,h_{k-l-j+1}\big)(y)$
$\qquad+O(\eta^{K+1})$
@@
Lemma A.7 (Bias and Variance of Gaussian Smoothing).
@@
$\left|\mathop{\mathbb{E}}[\widehat{q}^{(m)}_{b}(y)]-q^{(m)}(y)\right|$ $=\left|\int\frac{1}{b^{m+1}}\phi^{(m)}\left(\frac{y-x}{b}\right)q(x)\,\mathrm{d}x-q^{(m)}(y)\right|$
$=\left|\int\frac{1}{b^{m}}\phi^{(m)}(z)q(y-bz)\,\mathrm{d}z-q^{(m)}(y)\right|$
$=\left|\int\frac{(-1)^{m}}{b^{m}}\phi(z)\frac{\mathrm{d}^{m}}{\mathrm{d}z^{m}}q(y-bz)\,\mathrm{d}z-q^{(m)}(y)\right|$
$=\left|\int\phi(z)q^{(m)}(y-bz)\,\mathrm{d}z-\int\phi(z)q^{(m)}(y)\,\mathrm{d}z\right|$
$=\left|\int\phi(z)\left(q^{(m)}(y-bz)-q^{(m)}(y)+q^{(m+1)}(y)bz\right)\,\mathrm{d}z\right|$
$\leq\int\phi(z)\left|q^{(m)}(y-bz)-q^{(m)}(y)+q^{(m+1)}(y)bz\right|\,\mathrm{d}z$
$\leq\int\phi(z)\frac{Lb^{2}z^{2}}{2}\,\mathrm{d}z=\frac{Lb^{2}}{2}\;,$
where the second last step uses the fact $\int z\phi(z)\,\mathrm{d}z=0$ and then by Taylor expansion of $q^{(m)}(y-bz)$ around $y$, and the last step uses the Hölder smoothness of $q$.
@@
$\mathrm{Var}\left[\frac{1}{b^{m+1}}\phi^{(m)}\left(\frac{y-Y_{1}}{b}\right)\right]$ $\leq\int\frac{1}{b^{2m+2}}\left(\phi^{(m)}\left(\frac{y-x}{b}\right)\right)^{2}q(x)\,\mathrm{d}x$
$=\int\frac{1}{b^{2m+1}}\left(\phi^{(m)}(z)\right)^{2}q(y-bz)\,\mathrm{d}z\;.$
@@
$\int\frac{1}{b^{2m+1}}\left(\phi^{(m)}(z)\right)^{2}q(y-bz)\,\mathrm{d}z$
$=\int\frac{1}{b^{2m+1}}He_{m}^{2}(z)\phi^{2}(z)q(y-bz)\,\mathrm{d}z$
$\leq\sup_{u,z}\left\{q(u-bz)\phi(z)\right\}\cdot\int\frac{1}{b^{2m+1}}He_{m}^{2}(z)\phi(z)\,\mathrm{d}z$
$\leq\frac{L}{\sqrt{2\pi}}\frac{1}{b^{2m+1}}\int He_{m}^{2}(z)\phi(z)\,\mathrm{d}z=\frac{Lm!}{\sqrt{2\pi}}\frac{1}{b^{2m+1}}\;,$
where the last step uses the fact $\int He_{m}^{2}(z)\phi(z)\,\mathrm{d}z=m!$. ∎
@@
Lemma A.8 (Higher-order Score Representation).
@@
$\mathop{\mathbb{E}}_{Y\sim Q}\frac{1}{2}\left|f(Y)-f^{\ast}_{m}(Y)\right|^{2}$ $=\mathop{\mathbb{E}}_{Y\sim Q}\left[\tfrac{1}{2}f(Y)^{2}+(-1)^{m+1}f^{(m)}(Y)\right]$
$\qquad-\mathop{\mathbb{E}}_{Y\sim Q}\left[\tfrac{1}{2}f^{\ast}_{m}(Y)^{2}+(-1)^{m+1}(f^{\ast}_{m})^{(m)}(Y)\right]\;.$
Proof of Lemma A.8.
@@
$\mathop{\mathbb{E}}_{Y\sim Q}\frac{1}{2}\left|f(Y)-f^{\ast}_{m}(Y)\right|^{2}$ $=\int\frac{1}{2}\left(f(y)-\frac{q^{(m)}(y)}{q(y)}\right)^{2}q(y)\,\mathrm{d}y$
$=\int\left[\frac{1}{2}f(y)^{2}q(y)-f(y)q^{(m)}(y)+\frac{1}{2}\frac{q^{(m)}(y)^{2}}{q(y)}\right]\,\mathrm{d}y$
$=\mathop{\mathbb{E}}_{Y\sim Q}\left[\frac{1}{2}f(Y)^{2}+(-1)^{m+1}f^{(m)}(Y)\right]+\int\frac{1}{2}\frac{q^{(m)}(y)^{2}}{q(y)}\,\mathrm{d}y\;.$
Here, the last step applies integration by parts $m$ times and uses the vanishing boundary conditions up to order m1m-1. Plugging in $f=f^{\ast}_{m}$, we have the higher-order Fisher-type information [Bob24] identity
$\int\frac{1}{2}\frac{q^{(m)}(y)^{2}}{q(y)}\,\mathrm{d}y$ $=-\mathop{\mathbb{E}}_{Y\sim Q}\left[\frac{1}{2}f^{\ast}_{m}(Y)^{2}+(-1)^{m+1}(f^{\ast}_{m})^{(m)}(Y)\right]\;.$